# Polynomial Lattices for the BIKE Cryptosystem

### v1

@@ Abstract
In this paper we introduce a rank 222 lattice over a polynomial ring arising from the public key of the BIKE cryptosystem [1]. The secret key is a sparse vector in this lattice. We study properties of this lattice and generalize the recovery of weak keys from [3]. In particular, we show that they implicitly solved a shortest vector problem in the lattice we constructed. Rather than finding only a shortest vector, we obtain a reduced basis of the lattice which makes it possible to check for more weak keys.
@@ 1 Introduction and Related Work
One of the round 444 candidates of the NIST competition for Post-Quantum cryptography was the BIKE cryptosystem [1]. It has not been selected for standardization, but the related scheme HQC [16] will be standardized.
@@ 2 The BIKE Cryptosystem and a Class of Weak Keys
First of all we describe the BIKE cryptosystem [1]. Let us fix a prime $r$ such that 222 generates $\mathbb{F}_{r}^{\times}$. The secret key consists of two sparse (weight $v\sim\sqrt{r}$) elements $h_{1},h_{2}\in\mathbb{F}_{2}[x]/(x^{r}-1)$. The public key is $\frac{h_{1}}{h_{2}}$. The goal of a key recovery attack is to recover $h_{1}$ and $h_{2}$ from $h$. Actually, we do not need the very same $h_{1}$ and $h_{2}$. Any two sparse elements $h_{1},h_{2}$ satisfying $h=\frac{h_{1}}{h_{2}}$ are good enough.
@@ 3 Lattices and Thue’s Lemma
Lemma 2.
Let $\gamma,b,a^{\ast},t^{\ast}\in\mathbb{Z}_{\geq 0}$. Assume $0<a^{\ast}<\gamma<a^{\ast}t^{\ast}$. Then there exist $a,t\in\mathbb{Z}$ such that $|a|<a^{\ast},0<|t|<t^{\ast}$ satisfying
@@
Under the assumption that the volume of a convex, symmetric (around 0) set $R$ is larger than $4\gamma$, we get, using Minkowski’s Theorem, a nonzero lattice point in $R$. The rectangle $R=\{(x_{1},x_{2}):|x_{1}|<a^{\ast},|x_{2}|<t^{\ast}\}$ has volume $4a^{\ast}t^{\ast}>4\gamma$. Hence there exists a nonzero lattice point in it. It is easy to see that if $t=0$ we also get $a=0$ and the statement is proven. ∎
[…]
@@ 4 Lattices over Polynomial Rings
For $f\in\mathbb{F}_{q}[x]$ we define $|f|=\deg f$ for $f\neq 0$ and |0|=|0|=$\infty$| 0 | = - ∞. We define the norm of $a=(a_{1},\ldots,a_{n})\in\mathbb{F}_{q}[x]^{n}$ as
@@
The first successive minimum is defined to be the norm $|m_{1}|$ of a shortest non-zero vector $m_{1}$ in the lattice $L$. Iteratively we define the $j$-th successive minimum for $2\leq j\leq n$ as follows. Take $j-1$ linearly independent vectors $m_{1},\ldots,m_{j-1}\in L$ that attain the corresponding successive minima. The $j$-th successive minimum is the norm $|m_{j}|$ of a shortest vector $m_{j}\in L$ that is linearly independent of $m_{1},\ldots,m_{j-1}$.
@@
We will prove the following two lemmata since we could not find the statements and the proofs in the literature. However, they are not strictly necessary for the remainder of the paper. It is a classical result that for lattices $L_{2}\subseteq L_{1}\subseteq\mathbb{R}^{n}$ we have $[L_{1}:L_{2}]=\frac{\operatorname{covol}(L_{2})}{\operatorname{covol}(L_{1})}$. We prove the following analogue.
@@
$[L_{1}:L_{2}]=q^{\left|\frac{\det(L_{2})}{\det(L_{1})}\right|}=q^{\operatorname{covol}(L_{2})-\operatorname{covol}(L_{1})}.$
@@
$[L_{1}:L_{2}]=\prod_{i=1}^{n}q^{|\alpha_{i}|}=q^{\sum_{i=1}^{n}|\alpha_{i}|}.$
@@
$[\mathbb{F}_{q}[x]^{n}:\Lambda_{g}^{\perp}(A)]\leq q^{|g|k}.$
@@ 5 BIKE from a Lattice Viewpoint
Note that this lattice can be constructed solely from the public key. We will propose a new algorithm to get weak keys in a similar way as in [3] by using lattice reduction and adding a possibility for using brute force. In practice our algorithm improves only very slightly upon the algorithm proposed before, but it gives a more unified perspective on it, especially since lattice reduction for polynomial rings over finite field is computationally easy in general and not just for the 222-dimensional case as for the integers.
@@
[…]
@@
[…]
@@
Let $L=\{(h_{1},h_{2}):h_{1}-hh_{2}\equiv 0\mod x^{r}-1\}$. Then one can find the successive minima of $L$ in time $O(r^{2})$.
@@
How hard is it to find short vectors in the lattice $L$ where the ”norm” is defined by $|(h_{1},h_{2})|=\operatorname{wt}(h_{1})+\operatorname{wt}(h_{2})$.
However, this is not even a norm and the natural norm on the lattice is quite far from the norm in Problem 20. Are there better algorithms to tackle this problem?
@@ 6 Conclusion and Open Questions
One of the main open questions is whether there are better key recovery attacks that use a similar approach. In particular, the lattice constructed in this paper is a rank 222 lattice and closely related to the Euclidean algorithm and continued fractions. Is there a way to construct a higher rank lattice that makes it possible to get (many) more keys?
We would like to point out that there have been two problems that started out with rank 222 lattices (or extended Euclidean algorithm and continued fractions) and could be improved considerably using higher rank lattices. One is Wiener’s classical attack [22] on small private keys using continued fractions and its improvements by Boneh and Durfee [8] relying on Coppersmiths attack [10] which uses the LLL algorithm. Another one is the decoding of Reed-Solomon codes by Berlekamp and Massey [5, 15] (reinterpreted as continued fractions by Mills [17], Welch and Scholtz [21]) and the breakthrough of List Decoding by Guruswami and Sudan [12] (reinterpreted by Bernstein [6], Cohn and Heninger [9] as an analogue of Coppersmith’s theorem). Can something similar be done for the key recovery of the secret key of BIKE?
@@ Acknowledgements
I would like to thank Maxime Bombar for introducing me to the Function Field Approach for quasi-cyclic codes, which underlies cryptosystems like BIKE or HQC. Furthermore, I would like to thank him for mentioning the work of Bernstein, Cohn and Heninger on the Coppersmith Lattice Approach to List Decoding. In addition I would like to thank Paolo Santini for bringing [3] to my attention.

### v2

@@ Abstract
In this paper we introduce a rank $2$ lattice over a polynomial ring arising from the public key of the BIKE cryptosystem [1]. The secret key is a sparse vector in this lattice. We study properties of this lattice and generalize the recovery of weak keys from [3]. In particular, we show that they implicitly solved a shortest vector problem in the lattice we constructed. Rather than finding only a shortest vector, we obtain a reduced basis of the lattice which makes it possible to check for more weak keys.
@@ 1 Introduction and Related Work
One of the round $4$ candidates of the NIST competition for Post-Quantum cryptography was the BIKE cryptosystem [1]. It has not been selected for standardization, but the related scheme HQC [11] will be standardized.
@@ 2 The BIKE Cryptosystem and a Class of Weak Keys
First of all we describe the BIKE cryptosystem [1]. Let us fix a prime $r$ such that $2$ generates $\mathbb{F}_{r}^{\times}$. The secret key consists of two sparse (weight $v\sim\sqrt{r}$) elements $h_{1},h_{2}\in\mathbb{F}_{2}[x]/(x^{r}-1)$. The public key is $\frac{h_{1}}{h_{2}}$. The goal of a key recovery attack is to recover $h_{1}$ and $h_{2}$ from $h$. Actually, we do not need the very same $h_{1}$ and $h_{2}$. Any two sparse elements $h_{1},h_{2}$ satisfying $h=\frac{h_{1}}{h_{2}}$ are good enough.
@@ 3 Lattices and Thue’s Lemma
Lemma 2 (Thue).
Let $\gamma,b,a^{\ast},t^{\ast}\in\mathbb{Z}_{\geq 0}$. Assume $0<a^{\ast}\leq\gamma<a^{\ast}t^{\ast}$. Then there exist $a,t\in\mathbb{Z}$ such that $|a|<a^{\ast},0<|t|<t^{\ast}$ satisfying
@@
Under the assumption that the volume of a convex, symmetric (around 0) body $R$ is larger than $4\gamma$, we get, using Minkowski’s Theorem, a nonzero lattice point in $R$. The rectangle $R=\{(x_{1},x_{2}):|x_{1}|<a^{\ast},|x_{2}|<t^{\ast}\}$ has volume $4a^{\ast}t^{\ast}>4\gamma$. Hence there exists a nonzero lattice point in it. It is easy to see that if $t=0$ we also get $a=0$ and the statement is proven. ∎
Note that we could also additionally assume $a\geq 0$ since we will get a nonzero lattice point $x$ and its inverse x-x. In [20] the author gives an efficient algorithm relying on the extended Euclidean algorithm to give an explicit solution for Lemma 2. Note that this amounts to finding a short vector in a rank $2$ lattice. So instead of the extended Euclidean algorithm we could run a lattice reduction algorithm like the Lagrange algorithm for rank $2$ lattices.
@@ 4 Lattices over Polynomial Rings
For $f\in\mathbb{F}_{q}[x]$ we define $|f|=\deg f$ for $f\neq 0$ and $|0|=$-$\infty$. We define the norm of $a=(a_{1},\ldots,a_{n})\in\mathbb{F}_{q}[x]^{n}$ as
@@
The first successive minimum is defined to be the norm $|m_{1}|$ of a shortest non-zero vector $m_{1}$ in the lattice $L$. Iteratively we define the $j$-th successive minimum for $2\leq j\leq n$ as follows. Take j1j-1 linearly independent vectors $m_{1},\ldots,m_{j-1}\in L$ that attain the corresponding successive minima. The $j$-th successive minimum is the norm $|m_{j}|$ of a shortest vector $m_{j}\in L$ that is linearly independent of $m_{1},\ldots,m_{j-1}$.
@@
We will prove the following two lemmata since we could not find the statements and the proofs in the literature. However, they are not strictly necessary for the remainder of the paper. It is a classical result that for lattices $L_{2}\subseteq L_{1}\subseteq\mathbb{R}^{n}$ we have $L_{1}:L_{2}]=\frac{\operatorname{covol}(L_{2})}{\operatorname{covol}(L_{1})}$. We prove the following analogue.
@@
$L_{1}:L_{2}]=q^{\left|\frac{\det(L_{2})}{\det(L_{1})}\right|}=q^{\operatorname{covol}(L_{2})-\operatorname{covol}(L_{1})}.$
@@
$L_{1}:L_{2}]=\prod_{i=1}^{n}q^{|\alpha_{i}|}=q^{\sum_{i=1}^{n}|\alpha_{i}|}.$
@@
$\mathbb{F}_{q}[x]^{n}:\Lambda_{g}^{\perp}(A)]\leq q^{|g|k}.$
@@ 5 BIKE from a Lattice Viewpoint
Note that this lattice can be constructed solely from the public key. We will propose a new algorithm to get weak keys in a similar way as in [3] by using lattice reduction and adding a possibility for using brute force. In practice our algorithm improves only very slightly upon the algorithm proposed before, but it gives a more unified perspective on it, especially since lattice reduction for polynomial rings over finite fields is computationally easy in general and not just for the $2$-dimensional case as for the integers.
@@
and consequently we can express $(h_{1},h_{2})$ as a linear combination of $(x^{r}-1,0)$ and $(h,1)$ with coefficients in $\mathbb{F}_{q}[x]$. ∎
@@
Let $L=\{(h_{1},h_{2}):h_{1}-hh_{2}\equiv 0\mod x^{r}-1\}$. There exists $a,b\in L$ with $|a|+|b|=r$. In particular there exists $a\in L$ with $|a|\leq\frac{r-1}{2}$.
@@
Let $L=\{(h_{1},h_{2}):h_{1}-hh_{2}\equiv 0\mod x^{r}-1\}$. Then one can find a reduced basis and hence also the successive minima of $L$ in time $O(r^{2})$.
@@
How hard is it to find short vectors in the lattice $L$ where the norm is defined by $|(h_{1},h_{2})|=\operatorname{wt}(h_{1})+\operatorname{wt}(h_{2})$.
However, while being a norm over $\mathbb{F}_{q}$ endowed with the trivial absolute value, this is not a norm over $\mathbb{F}_{q}(T)$ endowed with the absolute value arising from the degree and the natural norm on the lattice is quite far from the norm in Problem 20. Are there better algorithms to tackle this problem?
@@ 6 Conclusion and Open Questions
One of the main open questions is whether there are better key recovery attacks that use a similar approach. In particular, the lattice constructed in this paper is a rank $2$ lattice and closely related to the Euclidean algorithm and continued fractions. Is there a way to construct a higher rank lattice that makes it possible to get (many) more keys?
We would like to point out that there have been two problems that started out with rank $2$ lattices (or the extended Euclidean algorithm and continued fractions) and could be improved considerably using higher rank lattices. One is Wiener’s classical attack [22] on small private keys using continued fractions and its improvements by Boneh and Durfee [8] relying on Coppersmith’s attack [10] which uses the LLL algorithm. Another one is the decoding of Reed-Solomon codes by Berlekamp and Massey [5, 16] (reinterpreted as continued fractions by Mills [17], Welch and Scholtz [21]) and the breakthrough of List Decoding by Guruswami and Sudan [13] (reinterpreted by Bernstein [6], Cohn and Heninger [9] as an analogue of Coppersmith’s theorem). Can something similar be done for the key recovery of the secret key of BIKE? At present this question is open and it requires new ideas.
@@ Acknowledgements
I would like to thank Maxime Bombar for introducing me to the Function Field Approach for quasi-cyclic codes, which underlies cryptosystems like BIKE or HQC. Furthermore, I would like to thank him for mentioning the work of Bernstein, Cohn and Heninger on the Coppersmith Lattice Approach to List Decoding. In addition I would like to thank Paolo Santini for bringing [3] to my attention. I would also like to thank the reviewers for their valuable comments.